OpenStudy (anonymous):

How do I solve (d+1/2)^2=1

OpenStudy (blurbendy):

d + / 2 ???

OpenStudy (blurbendy):

oh d + 1 all over 2?

OpenStudy (anonymous):

if \((d+\frac{1}{2})^2=1\) then \(d+\frac{1}{2}=1\) or \(d+\frac{1}{2}=-1\)

OpenStudy (anonymous):

Im trying to find the roots by factoring